OpenStudy (anonymous):

The factorial 4! is equal to

OpenStudy (mimi_x3):

\[4!=4*3*2*1\]

OpenStudy (anonymous):

24?

OpenStudy (mimi_x3):

Yes!

OpenStudy (unklerhaukus):

good stuff

OpenStudy (mimi_x3):

Or just use a calculator that has the factorial sign

OpenStudy (lgbasallote):

i tried finding the factorial of LGBAAA!!! in my scientific calculator. said "syntax error"

OpenStudy (anonymous):

\[\large n! = n \times (n-1) \times (n-2) \times (n-3) \times .............. 3 \times 2 \times 1\] Put n= 4 here..

OpenStudy (anonymous):

@waterineyes n! = 4 * 3 * 2 * 1 * 3 * 2 * 1 = 144 O:

OpenStudy (lgbasallote):

lol should only be \[n! = n \times (n-1) \times (n-2) \times \cdots \times 1\]

OpenStudy (mimi_x3):

its the same thing igba

OpenStudy (mimi_x3):

Unrolling factorials.

OpenStudy (lgbasallote):

nope... x 3 x 2 x 1 is different like @nphuongsun93 said

OpenStudy (anonymous):

Really??? Well, I got: \[4! = 4 \times (4-1) \times (4-2) \times (4-3) = 4 \times 3 \times 2 \times 1 = 24\]

OpenStudy (lgbasallote):

where's x 3 x 2 x 1 there? lol

OpenStudy (mimi_x3):

What waterineyes wrote is the same as what Igba wrote; but igba just wrote it in a shorter form. From memory you can also wrote it as: \(n! = n(n-1)!\)

OpenStudy (anonymous):

@waterineyes i'm just teasing you we all know this

OpenStudy (anonymous):

That in the decreasing sense I wrote there... You are also right lgba.. if we put 1 in the last then this will reduce confusion..

OpenStudy (lgbasallote):

but the dotdotdot does express something ;)

OpenStudy (anonymous):

No @Mimi_x3 you can also right it as: \[n! = n(n-1)(n-2)!\] Ha ha ha..

OpenStudy (anonymous):

No you are wrong lgba.. Careful..

OpenStudy (lgbasallote):

?

OpenStudy (anonymous):

\[n! = (n(n-n)!)!\]

OpenStudy (mimi_x3):

Wel, from my notes \(n! = (n-1)!*n\) Mynotes cannot be wrong.

OpenStudy (anonymous):

@Mimi_x3 you can subtract 1 upto the length you want..

OpenStudy (lgbasallote):

\[n! = (n \times 1)!\]

OpenStudy (anonymous):

\[n! = (n \times 0!)!\]

OpenStudy (anonymous):

\[ \huge \huge n! = n! \]

OpenStudy (lgbasallote):

\[n! = 42!\]

OpenStudy (anonymous):

\[\large n! = \sqrt{(n!)^2}\]